Starting from home, you walked 200.0 m N going to the market. From there, you proceeded 300.0 m E to a grocery store before walking 250.0 m SE to buy computer parts. You decided then to go to school which is 100.0 m 30 degrees N of E of the computer shop. Applying the concepts of vector addition using the graphical and component method, what is your latest location with respect to your house? Show complete solutions of both methods used.


1
Expert's answer
2021-11-04T10:17:11-0400

"y=200-250\cdot\cos45°+100\cdot\sin30°=73.22\ (m)"


"x=300+250\cdot\sin45°+100\cdot\cos30°=536.4\ (m)"


"r=\sqrt{73.22^2+536.4^2}=568.1\ (m)"